Evaluate the definite integral

                                                    

                    300

∫            ( 2,1 + 7 / 1000  T ) dT

                     150

 


a.551


b.300


c.150


d.450



1
Expert's answer
2022-05-03T16:23:13-0400
"\displaystyle\int_{150}^{300}(2.1+\dfrac{7}{1000}T)dT=[2.1T+0.0035T^2]\begin{matrix}
 300 \\
150
\end{matrix}"

"=551.25"

a.551